North Corp. EBIT is $200. It has a debt-equity ratio of 25% and a WACC of 16%. Debt

interest is 12%. Without taxes, what is the value of the equity? What is the value of the

firm? What is the cost of equity capital?


1
Expert's answer
2021-05-09T13:16:09-0400

we will find it using the formulas:

"DFL=\frac{EBIT}{EBIT-Interests}"

DFL=0.25

"0.25=\frac{200}{200-Interests}"

Interests=-600


"I=debt\times r"

"debt=\frac{I}{r}=\frac{-600}{0.12}=-5000"


"debt-equity ratio=\frac{L}{E}"

L=5000

"0.25=\frac{5000}{E}"

E=20 000


"value of the firm=E+L=20 000+5000=25000"


"WACC=Ks\times Ws+Kd\times Wd"

"0.16=Ks\times \frac{20 000}{25 000}+0.12\times \frac{5000}{25000}"

"0.16=Ks\times 0.8+0.12\times0.2"

"0.136=Ks\times 0.8"

Ks=0.17 or 17%